# The OATMEAL Survey. II. The 3D spin-orbit obliquity of an eccentric transiting brown dwarf in the Ruprecht 147 open cluster

### v1

@@ 3.1 The 3D spin-orbit obliquity
With our newly measured projected obliquity in hand, we may determine the 3D spin-orbit obliquity of the system. This is the angle between the spin axis of the primary star and the orbital axis of the brown dwarf companion. by combining $\lambda$ with the stellar spin axis inclination angle, $i_{\star}$, and the orbital inclination angle of the BD, $i_{0}$. The 3D obliquity may be calculated via:
@@
This form of the relationship between $\psi$, $\lambda$, $i_{\star}$, and $i_{0}$ follow from S. H. Albrecht et al. (2022). The first report of the rotation period of CWW 89A comes from G. Nowak et al. (2017) on an analysis of the K2 light curve showing a $P_{\rm rot}=12.6\pm 2.1$ days. If we adopt this value and the $v\sin{i_{\star}}=4.1\pm 0.4\,(1.0)\,{\rm km\,s^{-1}}$ from G. Nowak et al. (2017), we may derive the stellar inclination as $i_{\star}=74.7^{+9.4^{\circ}}_{-18.6}$ (Figure 2). The parenthetical uncertainty on $v\sin{i}$ is the value we adopt given the typical challenges in small $v\sin{i}$ measurements; our RM analysis also provides constraints on $v\sin{i}$ (Table 3) and both $v\sin{i}$ values are consistent with each other. This derivation follows that first outlined in K. Masuda & J. N. Winn (2020) and improved on in B. P. Bowler et al. (2023) using a bootstrap analysis of the relationship:
@@
Combining the results from this and Equation 1 yields a 3D obliquity value of $\psi=15.1^{+15.0^{\circ}}_{-10.9}$ for CWW 89Ab (Figure 3). This can be compared to $\psi$ measured for other transiting BDs in Table 2, making CWW 89Ab the third transiting BD with a measurement of $\psi$.
Figure 3: The posterior distribution of $\psi$ from sampling in $\cos{\psi}$ parameter space. Given the skewed nature of this distribution, a more conservative estimate of the 3D obliquity is $\psi\leq 30.1^{\circ}$ with the peak of the distribution at $\psi_{\rm peak}=3.1^{\circ}$.
@@ 4.1 The CWW 89 system
A collection of past studies (J. L. Curtis et al., 2013b; G. Torres et al., 2018; G. Nowak et al., 2017; T. G. Beatty et al., 2018; T. W. Carmichael et al., 2019) have helped construct a nearly-complete picture for the formation and evolution of the brown dwarf in the CWW 89 system. This section serves to contextualize the present work into this picture with the aim of further completing it with our new measurements of the orbital architecture of CWW 89. Here are the following key features of the CWW 89 system:
  • CWW 89A is a $1.0\,M_{\sun}$ star that hosts a distant $0.5\,M_{\sun}$ M-dwarf companion, CWW 89B, with a projected separation of 25 AU, or 81 mas, and $\Delta K=2.2$ dimmer than CWW 89A (T. G. Beatty et al., 2018)
  • CWW 89 is a probable member of the Ruprecht 147 open cluster (J. L. Curtis et al., 2013b) with an age of $2.76\pm 0.61$ Gyr (G. Torres et al., 2018)
  • The brown dwarf, CWW 89b, may have formed via core accretion given evidence presented in T. G. Beatty et al. (2018). In summary, secondary eclipses from Spitzer/IRAC (G. G. Fazio et al., 2004) indicate an over-luminous atmosphere of CWW 89Ab that is best explained by a temperature inversion caused by a superstellar C/O ratio in the brown dwarf. This favors a core accretion formation scenario.
  • The tidal quality factor of CWW 89b is estimated to have a lower limit of $Q_{b}\geq 10^{4.15}$ (T. G. Beatty et al., 2018), which yields a circularization timescale $\tau_{\rm circ}>7.5$ Gyr (T. W. Carmichael et al., 2019)
Currently, it is unclear whether or not the secondary star, CWW 89B, is coplanar or not with CWW 89Ab—an especially important aspect to consider given the implications for the M-dwarf’s influence on the orbital evolution of the brown dwarf.
@@ 4.2 Evidence for primordial alignment
Several factors affect the spin-orbit histories and outcomes for giant planet and brown dwarf companions to stars. Three key factors are: mass ratio $q$, scaled semi-major axis $a/R_{\star}$, and host star effective temperature. We show Figure 4 to illustrate the observed distribution in spin-orbit alignment $\lambda$ as a function of $T_{\rm eff}$ or $q$ (adapted from a similar figure in J. Rusznak et al. (2024)). These factors strongly dictate the circularization timescale $\tau_{\rm circ}$ and realignment timescale $\tau_{\rm CE}$ for giant planets and brown dwarfs, but importantly, these effects appear to be distinct between the two populations. From S. Albrecht et al. (2012), the realignment timescale is:
@@
for convective outer envelope (hence the CE subscript) stars $T_{\rm eff}\lesssim 6250$K. We derive $q=0.036$ and $a/R_{\star}=12.608$ for CWW 89Ab.
From this, it follows that less-massive companions (smaller mass ratio $q$) would have longer realignment timescales than more-massive companions around cool stars. The same conclusion is drawn for companions in larger scaled semi-major axis configurations. In the case of CWW 89A, Equation 3 indicates the realignment timescale is $\tau_{\rm CE}\approx 10$ Gyr, and greater than the $2.76\pm 0.61$ Gyr age of the system by $>$$3\sigma$. This implies that the spin-orbit alignment of the brown dwarf is primordial. On the contrary, when considering only the primary star, the provenance of the eccentricity of the brown dwarf is ambiguous given T. W. Carmichael et al. (2019) report the circularization timescale $\tau_{\rm circ}>7.5$ Gyr using the tidal quality factor of $Q_{b}\gtrsim 10^{4.15}$ for the brown dwarf from T. G. Beatty et al. (2018). This is again much longer than the system’s age, but the fact that the eccentricity is $e=0.19$ means that this can be primordial or a result of interactions with the outer companion, CWW 89B. T. G. Beatty et al. (2018) report the $Q_{b}$ based on an extrapolation using rotational angular momentum evolution calculations between a star and giant planet from J. Leconte et al. (2010). This calculation attempts to account for the exchange in energy via tides that result in the spin-up of the host star rotation rate and the increase in internal heat of the companion.
@@ 4.2.1 Coplanar high-eccentricity migration
This brings us to one theory that could explain the non-zero eccentricity of CWW 89Ab while preserving a zero (aligned) spin-orbit angle. C. Petrovich (2015) present a framework that plausibly describes the story of CWW 89 via a coplanar high-eccentricity migration (coplanar HEM) for the BD driven by the outer M-dwarf. In this scenario, a distant companion could be responsible for guiding the BD inwards to its current configuration should the following conditions be satisfied:
@@
where $m$ and $a$ are the respective initial masses and semi-major axes of the inner (CWW 89Ab; $a_{\rm in}=0.06$ AU, $m_{\rm in}=M_{b}=39\,M_{\rm J}=0.04\,M_{\sun}$) and outer companion (CWW 89B; $a_{\rm out}\gtrsim 25$ AU, $m_{\rm out}=M_{B}=0.5\,M_{\sun}$). In the case of CWW 89, we are limited with a projected separation for CWW 89B, but even in the most conservative scenario (i.e. the minimum allowable $a_{\rm out}$) for this framework, the above inequality still holds, meaning that CWW 89B may have induced coplanar high-eccentricity migration of CWW 89Ab. This inequality assumes an initial eccentricity of the inner companion $e\gtrsim 0.5$, so we would require the eccentricity of the inner brown dwarf to sufficiently dampen from $e\approx 0.5$ to $e=0.2$ for its present day configuration.
Figure 4: Top: $\lambda$ vs $T_{\rm eff}$ for transiting companions $\gtrsim 0.3\,M_{\rm J}$. The Kraft Break regime in the vertical shaded region is from A. C. Beyer & R. J. White (2024), with the traditional 6250K threshold denoted by the dashed line. Brown dwarfs are represented by square symbols. CWW 89Ab is the lowest-mass transiting brown dwarf around a sub-Kraft Break star with its $\lambda$ and $\psi$ measured. Bottom: $\lambda$ vs $q$ ($M_{b}/M_{\star}$) for the same sample in the top panel. We highlight the threshold ($q=2\times 10^{-3}$) past which primordial alignment ($\lambda\approx 0^{\circ}$) dominates the observed distribution (credit to J. Rusznak et al., 2024). Circled points are systems with inner transiting brown dwarfs and confirmed outer companions (CWW 89 and LP 261-75) or Gaia RUWE values above 1.5. All of the inner brown dwarfs except for the one in LP 261-75 are on eccentric orbits ($e>0.18$), plausibly following a coplanar HEM scenario. The misaligned giant planet system at $q\approx 10^{-2}$ is XO-3b discussed in Section 4.2.2. Data in both panels are from Figure 2 of J. Rusznak et al. (2024).
@@ 4.2.2 Other massive hierarchical triple systems
There are 4 other transiting systems that feature an inner massive ($\gtrsim 10\,M_{\rm J}$) companion with a measured spin-orbit angle and with confirmed or plausibly outer stellar companions. These are TOI-2119 (C. I. Cañas et al., 2022; T. W. Carmichael et al., 2022; L. Doyle et al., 2025), and TOI-2533 (S. P. Schmidt et al., 2023; T. Ferreira dos Santos et al., 2024), LP 261-75 (J. M. Irwin et al., 2018; M. Brady et al., 2025), and XO-3 (J. N. Winn et al., 2008; J. Rusznak et al., 2024). Given how suitably the coplanar high-eccentricity migration scenario could explain the CWW 89 system, we examine how it might apply to these other systems.
@@
    What is absent from our current knowledge of these other 2 more-massive BDs is a confirmation of an outer stellar companion like the detection of CWW 89B. L. Doyle et al. (2025) cite the relatively large re-normalized unweighted error (RUWE, V. Belokurov et al., 2020; Gaia Collaboration et al., 2023) from the Gaia mission of 1.93 for TOI-2119 as an indicator of an unresolved massive companion. An RUWE $>$ 1.4 indicates the strong potential for a binary system (V. Belokurov et al., 2020; D. M. Krolikowski et al., 2021).
@@
    Notably, TOI-2119 is an M-dwarf and could join other M-dwarf primary hierarchical triple systems with an inner transiting BD and outer secondary M-dwarf if the high RUWE is proof of an outer companion. These other hierarchical systems are: LHS 6343 (J. A. Johnson et al., 2011), NLTT 41135 (J. Irwin et al., 2010), NGTS-7 (J. A. G. Jackman et al., 2019), NGTS-28 (B. A. Henderson et al., 2024b), and LP 261-75 (J. M. Irwin et al., 2018; M. Brady et al., 2025). These 5 systems are a part of the total 9 M-dwarf primaries known to host transiting BDs ($\sim 56\%$); the other 4, TOI-2119 included, do not have confirmed outer companions. This emerging trend appears to show BDs that transit M-dwarfs to have a reasonable chance to be part of a hierarchical system with an additional outer M-dwarf. We note that one recent study by C. Cifuentes et al. (2025) indicates that the M-dwarf multiplicity fraction for separations of 0.01 to $10^{5}$ AU ranges from 26% to 42% depending on the confirmation of candidate stellar companions. However, we do robustly not confirm that this is a trend for M-dwarf hosts to transiting BDs here, so the best quantitative evidence for TOI-2119 to host an unresolved M-dwarf companion is its inflated RUWE.
@@
    The TOI-2533 system has an RUWE of 1.57, but is an F8 star. With only circumstantial evidence for the existence of an outer companion in the TOI-2119 and TOI-2533 systems via the Gaia RUWE exceeding a value of 1.4, it is not clear if an outer companion would extend the shared characteristics between TOI-2119, TOI-2533, and CWW 89, thus fitting the same coplanar HEM scenario from C. Petrovich (2015). Lastly, if TOI-2119 and TOI-2533 host a secondary M-dwarf companion, we could conclude that the condition in Equation 4 would be satisfied, since we could assume the hypothetical companion orbits exterior to the brown dwarf.
@@
  • Misaligned inner super-Jupiter companion: XO-3b is a massive $12\,M_{\rm J}$ planet in a misaligned ($\lambda=37.8\pm 3.7.^{\circ}$) and eccentric ($e=0.28$) 3-day orbit (J. N. Winn et al., 2008). This makes XO-3 the singular exception to the observed trend in aligned systems with high mass ratio (Figure 4). J. Rusznak et al. (2024) are the first to point out the RUWE of 1.25 for this system and later confirm the presence of a stellar companion via high-resolution imaging at 90AU. J. Rusznak et al. (2024) argue that this outer stellar companion may have been more of a pivotal factor in the misalignment of XO-3b than the effective temperature of the primary star being above the Kraft Break. This system is the only outlier among other massive companions ($M_{b}\gtrsim 10\,M_{\rm J}$) with well-determined $\lambda$ and it is unclear if the outer companion plays a role in this in the same fashion the outer companion to CWW 89A does.
@@
J. Rusznak et al. (2024) argue that the increase in planet-to-star mass ratio correlates with a preference for aligned spin-orbit angles for single-star systems. Their conclusions are only valid for single-star systems because these are the objects that they limited their sample to. They provide an explanation for this spin-orbit alignment phenomenon in the circumstellar disks from which these giant planets and brown dwarfs grew. For those systems that experience dynamically disruptive events early in their formation process, especially in low-mass disks with less mass available to accrete, they may experience a kick that can either or both drive an increase in eccentricity (i.e. high-eccentricity migration) and a misalignment with the plane of the disk and host star, thus depriving the planet of any additional material to grow in mass.
@@
This argument is presented to explain the ubiquity of misaligned sub-Saturn planets around stars above and below the Kraft Break. The opposite of this is seen for brown dwarf-mass and even super-Jupiter mass (i.e. $\gtrsim 2\,M_{\rm J}$) where every system is aligned around hot and cool stars (Figure 4). It is then the Jupiter-mass planets which represent a transitional region in mass where the effects of early-time dynamical excitation affects these planets less strongly than sub-Saturns, but more strongly than super-Jupiters and brown dwarfs. This is a summary of the argument established in J. Rusznak et al. (2024) and it is clear that the transiting brown dwarf population follows this very closely; CWW 89Ab being no exception save for the fact that it is not a single-star system as J. Rusznak et al. (2024) require for their argument to be valid (allowing them to exclude the XO-3 system).
@@
Of the super-Jupiter and brown dwarf systems with measured spin-orbit angles and RUWE $\gtrsim 1.25$, only 1 (XO-3b) in 4 total (CWW 89, TOI-2119, and TOI-2533) is misaligned. As this particular sample continues to grow, we expect it to reveal how the presence of outer companions to transiting brown dwarfs affect spin-orbit alignment. This work and the recent few confirming the aligned orbits of transiting brown dwarfs with the host star’s spin axis all but confirms that the underlying obliquity distribution for these massive companions is consistent with $\lambda\approx 0^{\circ}$ as G. Hébrard et al. (2011) establish early in the era of spin-orbit measurements for giant planet and subsequent works by K. Hixenbaugh et al. (2023); T. Gan et al. (2024); S. Giacalone et al. (2024) reinforce.

### v2

@@ 3.1 The 3D spin-orbit obliquity
With our newly measured projected obliquity in hand, we may determine the 3D spin-orbit obliquity of the system. This is the angle between the spin axis of the primary star and the orbital axis of the brown dwarf companion. By combining $\lambda$ with the stellar spin axis inclination angle, $i_{\star}$, and the orbital inclination angle of the BD, $i_{0}$, the 3D obliquity may be calculated via:
@@
This form of the relationship between $\psi$, $\lambda$, $i_{\star}$, and $i_{0}$ follows from albrecht2022. The first reported rotation period of CWW 89A comes from nowak17 on an analysis of the K2 light curve showing a $P_{\rm rot}=12.6\pm 2.1$ days. If we adopt this value and the $v\sin{i_{\star}}=4.1\pm 0.4\,(1.0)\,{\rm km\,s^{-1}}$ from nowak17, we may derive the stellar inclination as $i_{\star}=74.4^{+9.7^{\circ}}_{-18.3}$ (Figure 3). The parenthetical uncertainty on $v\sin{i}$ is the value we adopt given the typical challenges in small $v\sin{i}$ measurements; our RM analysis also provides constraints on $v\sin{i}$ (Table 4) and both $v\sin{i}$ values are consistent with each other. This derivation follows that first outlined in masuda2020 and improved on in bowler2023_inclination using a bootstrap analysis of the relationship:
@@
An artifact of the closeness in value of $v\sin{i_{\star}}$ and the rotational velocity ($v_{\rm rot}=2\pi R_{\star}/P_{\rm rot}$) is a broad distribution in $\cos{i_{\star}}$; this occurs because the argument of Equation 2 must be <1<1, or $v\sin{i_{\star}}<v_{\rm rot}$. This is illustratively explained best in Figure 1, top left panel of masuda2020. In any case, combining the results shown in Figure 3 here and Equation 1 yields a 3D obliquity value of $\psi=15.1^{+15.0^{\circ}}_{-10.9}$ for CWW 89Ab (Figure 4). This can be compared to $\psi$ measured for other transiting BDs in Table 3, making CWW 89Ab the third transiting BD with a measurement of $\psi$. We note that our measurements of $\lambda$ and $\psi$ are consistent within 1-$\sigma$ of concurrent measurements taken by zak2025.
Figure 4: The posterior distribution of $\psi$ from sampling in $\cos{\psi}$ parameter space. Given the skewed nature of this distribution (driven by $v\sin{i_{\star}}\lessapprox v_{\rm rot}$), a more conservative estimate of the 3D obliquity is $\psi\leq 30.1^{\circ}$ with the peak of the distribution at $\psi_{\rm peak}=3.1^{\circ}$.
@@ 4.1 The CWW 89 system
A collection of past studies (curtis2013; r147_age; nowak17; cww89a; carmichael19) have helped construct a nearly complete picture for the formation and evolution of the brown dwarf in the CWW 89 system. This section contextualizes the present work and this picture with the aim of further completing it with our new measurements of the spin-orbit obliquity. Summarized here are the key features of the CWW 89 system:
  • CWW 89A is a $1.0\rm\,M_{\sun}$ star that hosts a distant $0.5\rm\,M_{\sun}$ M-dwarf companion, CWW 89B, with a projected separation of 25 AU, or 81 mas, and $\Delta K=2.2$ dimmer than CWW 89A (cww89a)
  • CWW 89 is a probable member of the Ruprecht 147 open cluster (curtis2013) with an age of $2.76\pm 0.61$ Gyr (r147_age)
  • The brown dwarf, CWW 89b, may have formed via core accretion given evidence presented in cww89a; in summary: secondary eclipses from Spitzer/IRAC (SpitzerIRAC) indicate an over-luminous atmosphere of CWW 89Ab that is best explained by a temperature inversion caused by a superstellar C/O ratio in the brown dwarf—this C/O ratio favors a core accretion formation scenario
  • The tidal quality factor of CWW 89b is estimated to have a lower limit of $Q_{b}\geq 10^{4.15}$ (for details, see section 5 of cww89a), which yields a circularization timescale $\tau_{\rm circ}>7.5$ Gyr (carmichael19)
[…]
Further pulling CWW 89Ab away from a planet-like formation origin is the fact that stellar multiples are often found in hierarchical configurations not dissimilar from the CWW 89 system (e.g. tokovinin2008; duchene2013; moe2017), which implies that it is more similar to them. A study by schlaufman2018 also finds evidence that core accretion is challenged to form companions that are more massive than $\gtrsim 4\,M_{\rm J}$, especially around stars with super-solar metallicity like CWW 89A at $\rm[Fe/H]=+0.2$. These lines of evidence that disfavor a core accretion scenario for CWW 89Ab must be weighed against the findings of cww89a for a more complete picture.
Currently, it is unclear whether or not the secondary star, CWW 89B, is or is not coplanar with CWW 89Ab—an especially important aspect to consider given the implications for the M-dwarf’s influence on the orbital evolution of the brown dwarf. We discuss the implications of CWW 89B’s non-coplanarity with CWW 89Ab in more detail in Section 4.3.2.
@@ 4.2 Evidence for primordial alignment
Several factors affect the spin-orbit histories and outcomes for giant planet and brown dwarf companions to stars. Three key attributes are: mass ratio $q$, scaled semi-major axis $a/R_{\star}$, and host star effective temperature. We show Figure 5 to illustrate the observed distribution in spin-orbit alignment $\lambda$ as a function of $T_{\rm eff}$ or $q$ (adapted from a similar figure in rusznak2024). These factors strongly dictate the circularization timescale $\tau_{\rm circ}$ and realignment timescale $\tau_{\rm CE}$ for giant planets and brown dwarfs, but importantly, these effects appear to be distinct between the two populations.
From albrecht2012, the realignment timescale is:
@@
for convective outer envelope (hence the CE subscript) stars $T_{\rm eff}\lesssim 6250$K. We derive $q=0.035$ and $a/R_{\star}=12.21$ for CWW 89Ab.
From this, it follows that lower-mass companions (smaller mass ratio $q$) would have longer realignment timescales than higher-mass companions around cool stars. The same conclusion is drawn for companions at larger scaled semi-major axes. In the case of CWW 89A, Equation 3 indicates the realignment timescale is $\tau_{\rm CE}\approx 10$ Gyr, and greater than the $2.76\pm 0.61$ Gyr age of the system by $3\sigma$. This implies that the spin-orbit alignment of the brown dwarf is primordial.
However, the provenance of the eccentricity of the brown dwarf is ambiguous given carmichael19 report the circularization timescale $\tau_{\rm circ}>7.5$ Gyr using the tidal quality factor of $Q_{b}\gtrsim 10^{4.15}$ for the brown dwarf from cww89a. This reported $Q_{b}$ is based on an extrapolation using rotational angular momentum evolution calculations between a star and giant planet from leconte2010_tidalQ. This calculation attempts to account for the exchange in energy via tides that result in the spin-up of the host star rotation rate and the increase in internal heat of the companion. The fact that the circularization timescale is much longer than the system’s age while the BD maintains an eccentricity $e=0.19$ means that this is only possibly primordial. If the BD was not formed in an eccentric orbit, then its current eccentricity may be explained by interactions with the outer companion, CWW 89B.
@@ 4.2.1 Coplanar high-eccentricity migration
This brings us to one theory that could explain the non-zero eccentricity of CWW 89Ab while preserving a zero (aligned) spin-orbit angle. petrovich2015 presents a framework that plausibly describes the story of CWW 89 via a coplanar high-eccentricity migration (coplanar HEM, or CHEM) for the BD driven by the outer M-dwarf. In this scenario, a distant companion could be responsible for guiding the BD inwards to its current configuration without $\lambda$ deviating from zero should the following conditions be satisfied:
@@
where $m$ and $a$ are the respective initial masses and semi-major axes of the inner (CWW 89Ab; $a_{\rm in}=0.06$ AU, $m_{\rm in}=M_{b}=39\,M_{\rm J}=0.04\,M_{\sun}$) and outer companion (CWW 89B; $a_{\rm out}\gtrsim 25$ AU, $m_{\rm out}=M_{B}=0.5\,M_{\sun}$). In the case of CWW 89, we are limited with a projected separation for CWW 89B, but even in the most conservative scenario (the minimum allowable $a_{\rm out}$) for this framework, the above inequality still holds. This inequality assumes an initial eccentricity of the inner companion $e\gtrsim 0.5$, so this would require the eccentricity of the inner brown dwarf to sufficiently dampen from $e\approx 0.5$ to $e=0.2$ for its present day configuration. However, an additional criterion that petrovich2015 establishes is that if we assume that the initial $e_{\rm in}\approx 0.5$, then for the system to be stable, the initial $e_{\rm out}$ must be similar ($e_{\rm in}\approx e_{\rm out}\approx 0.5$) and the mass ratio between the inner and outer companion must be $m_{\rm in}/m_{\rm out}>0.36$ (see Figure 3 from petrovich2015). While the present and initially eccentricities of the outer companion are not known, the mass ratio between CWW 89Ab and CWW 89B is known. With an $m_{\rm in}/m_{\rm out}\approx 0.08$, the CWW 89 system falls short of the mass ratio criterion.
Given this, we consider two alternate scenarios: 1) An initially circular orbit for the BD that was subsequently excited into an eccentric orbit by the outer companion (discussed further in Section 4.3) or 2) an adherence to the assumption that the BD formed in an eccentric orbit, but in this case, the BD torqued the circumstellar disk due to its high mass. The mass of the BD taken with the evidence that it formed via core accretion (cww89a) may be interpreted as evidence that the circumstellar disk was eccentric and possibly torqued by the BD. Past works have confirmed that such disks have been observed to be eccentric (e.g. kley2006; romanova2023; romanova2024; commercon2024).
In both of these scenarios, the spin-orbit obliquity need not be misaligned during the inward high-eccentricity migration process of the BD. An example of this may be seen in the well-aligned ($\lambda=1.2\pm 2.8^{\circ}$) but eccentric ($e=0.72$) gas giant TOI-3362b (espinoza2023_CHEM). When taking these various factors into consideration, we are unable to decisively conclude whether or not CWW 89 is stable under the CHEM framework despite its cursory appearance of being an exemplary CHEM system like TOI-3362. Even so, the extreme mass and separation scales of CWW 89 (compared to those considered in petrovich2015) are not yet thoroughly tested in the CHEM framework and may not be completely ruled out.
Figure 5: Top: $\lambda$ vs $T_{\rm eff}$ for transiting companions $\gtrsim 0.3\,M_{\rm J}$. The Kraft Break regime in the vertical shaded region is from beyer_white2024, with the traditional 6250K threshold denoted by the dashed line. Brown dwarfs are represented by square symbols. CWW 89Ab is the lowest-mass transiting brown dwarf around a sub-Kraft Break star with its $\lambda$ and $\psi$ measured. Bottom: $\lambda$ vs $q$ ($M_{b}/M_{\star}$) for the same sample in the top panel. We highlight the threshold ($q=2\times 10^{-3}$) past which primordial alignment ($\lambda\approx 0^{\circ}$) dominates the observed distribution (credit to rusznak2024). Circled points are systems with inner transiting brown dwarfs and confirmed outer companions (CWW 89 and LP 261-75) or Gaia RUWE values above 1.5. Both CWW 89A and LP 261-75 have RUWE values <1.4<1.4 despite a confirmed stellar companion. All of the inner brown dwarfs except for the one in LP 261-75 are on eccentric orbits ($e>0.18$), plausibly following a coplanar HEM scenario. The misaligned giant planet system at $q\approx 10^{-2}$ is XO-3b discussed in Section 4.2.2. Data in both panels are from Figure 2 of rusznak2024.
@@ 4.2.2 Other massive hierarchical triple systems
There are 4 other transiting systems that feature an inner massive ($\gtrsim 10\,M_{\rm J}$) companion with a measured spin-orbit angle and with confirmed or plausible outer stellar companions. These are TOI-2119 (canas2021; carmichael2022; doyle2025), and TOI-2533 (Schmidt2023; dosSantos2024), LP 261-75 (irwin18; brady2025), and XO-3 (xo3b; rusznak2024). Given how suitably the coplanar high-eccentricity migration scenario could explain the CWW 89 system, we examine how it might apply to these other systems.
@@
    What is absent from our current knowledge of these other 2 BDs is a confirmation of an outer stellar companion like with the detection of CWW 89B. doyle2025 cite the relatively large re-normalized unweighted error (RUWE, ruweDR2; ruwedr3) from the Gaia mission of 1.93 for TOI-2119 as an indicator of an unresolved massive companion. An RUWE >> 1.4 indicates the strong potential for a binary system (ruweDR2; Krolikowski2021_RUWE).
@@
    Notably, TOI-2119 is an M-dwarf and could join other M-dwarf primary hierarchical triple systems with an inner transiting BD and outer secondary M-dwarf if the high RUWE is strong evidence of an outer companion. These other hierarchical systems are: LHS 6343 (johnson11_bd), NLTT 41135 (irwin10), NGTS-7 (jackman2019), NGTS-28 (ngts28A), LP 261-75 (irwin18; brady2025), and TOI-5389 (larsen2025). These 6 systems are a part of the total 14 M-dwarf primaries known to host transiting BDs ($\sim 43\%$); the other 8, TOI-2119 included, do not have confirmed outer companions. This emerging trend appears to show BDs that transit M-dwarfs to have a reasonable chance to be part of a hierarchical system with an additional outer M-dwarf. We note that one recent study by cifuentes2025 indicates that the M-dwarf multiplicity fraction for separations of 0.01 to $10^{5}$ AU ranges from 26% to 42% depending on the confirmation of candidate stellar companions. However, we do not robustly confirm that this is a trend for M-dwarf hosts to transiting BDs here, so the best quantitative evidence for TOI-2119 to host an unresolved M-dwarf companion is its inflated RUWE.
@@
    The TOI-2533 system has an RUWE of 1.57, but is an F8 star. With only circumstantial evidence for the existence of an outer companion in the TOI-2119 and TOI-2533 systems via the Gaia RUWE exceeding a value of 1.4, it is not clear if an outer companion would extend the shared characteristics between TOI-2119, TOI-2533, and CWW 89, thus fitting the same coplanar HEM scenario from petrovich2015. We point out that the RUWE for CWW 89A is only 1.2.
    A study by castro-ginard2024_RUWE demonstrated that the nominal RUWE binarity threshold of 1.4 can vary from $\sim$1.15 to $\sim$1.37 depending on the location in the sky (e.g. for crowded field calibrations) of the system due to the scanning angle of Gaia. There are also other proxies from Gaia for binarity are used in tandem with RUWE (e.g the astrometric acceleration; brandt2021; JZhang2024). Given this, we defer to a clear direct detection when arguing for the existence of an outer companion in this context.
    Lastly, if TOI-2119 and TOI-2533 host a secondary M-dwarf companion, we could conclude that the condition in Equation 4 would be satisfied, since we could assume the hypothetical companion orbits exterior to the brown dwarf. Given the much higher masses of TOI-2119b and TOI-2533b compared to CWW 89Ab, the mass ratio criterion from petrovich2015 would be easier to satisfy if the potential unseen companions in these 2 other systems were roughly $0.2\,M_{\sun}$.
@@
  • Misaligned inner super-Jupiter companion: XO-3b is a massive $12\,M_{\rm J}$ planet in a misaligned ($\lambda=37.8\pm 3.7.^{\circ}$) and eccentric ($e=0.28$) 3-day orbit (xo3b). This makes XO-3 the singular exception to the observed trend in aligned systems with high mass ratio (Figure 5). rusznak2024 are confirm the presence of a stellar companion via high-resolution imaging at 90AU for this system and highlight the RUWE of 1.25. rusznak2024 argue that this outer stellar companion may have been the dominant factor in the misalignment of XO-3b over the effective temperature of the primary star being above the Kraft Break. This system is the only outlier among other massive companions ($M_{b}\gtrsim 10\,M_{\rm J}$) with well-determined $\lambda$ and it is unclear if the outer companion plays a role in this in the same fashion the outer companion to CWW 89A does.
@@
rusznak2024 argue that the increase in planet-to-star mass ratio correlates with a preference for aligned spin-orbit angles for single-star systems. Their conclusions are intended to be valid for single-star systems because these are the objects that they limited their sample to. They provide an explanation for this spin-orbit alignment phenomenon in the circumstellar disks from which these giant planets and brown dwarfs grew. For those systems that experience dynamically disruptive events early in their formation process, especially in low-mass disks with less mass available to accrete, they may experience a kick that can either or both drive an increase in eccentricity (i.e. high-eccentricity migration) and a misalignment with the plane of the disk and host star, thus depriving the planet of any additional material to grow in mass.
@@
This argument is presented to explain the ubiquity of misaligned sub-Saturn planets around stars above and below the Kraft Break. The opposite of this is seen for brown dwarf-mass and even super-Jupiter mass (i.e. $\gtrsim 2\,M_{\rm J}$) companions where every system is aligned around hot and cool stars (Figure 5). It is then the Jupiter-mass planets which represent a transitional region in mass where the effects of early-time dynamical excitation affects these planets less strongly than sub-Saturns, but more strongly than super-Jupiters and brown dwarfs. This is a summary of the argument established in rusznak2024 and it is clear that the transiting brown dwarf population follows this very closely; CWW 89Ab being no exception save for the fact that it is not a single-star system as rusznak2024 require for their argument to be valid (allowing them to exclude the XO-3 system).
@@
Of the super-Jupiter and brown dwarf systems with measured spin-orbit angles and RUWE $\gtrsim 1.25$, only 1 (XO-3b) in 4 total (CWW 89, TOI-2119, and TOI-2533) is misaligned. As this particular sample continues to grow, we expect it to reveal how the presence of outer companions to transiting brown dwarfs affect spin-orbit alignment. This work and the recent few confirming the aligned orbits of transiting brown dwarfs with the host star’s spin axis may have growing implications at an underlying obliquity distribution for these massive companions is consistent with $\lambda\approx 0^{\circ}$. This would follow the conclusions that hebrard2011 establish early in the era of spin-orbit measurements for giant planet and subsequent works by hixenbaugh2023; gan2024; oatmeal1 would later reinforce.
Though the total sample for transiting brown dwarfs with measurement spin-orbit obliquities is quite small at 9 systems, the sample of transiting companions $\gtrsim 2\,M_{\rm J}$ is larger by a factor of 3. For this sample, the evidence is stronger for a $\lambda\approx 0^{\circ}$ spin-orbit obliquity and future investigations may find this parameter space of $\lambda$ vs. $q$ appropriate for relating the super-Jupiter and brown dwarf populations to one another.